# Sayantan Roy Sankha S. Basu Mihir K. Chakraborty School of Cognitive Science

### v1

@@ 4 Suszko Reduction: A Generalised Perspective
    Given an $S$-semantics $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$, the type-I $S$-entailment relation on $\mathscr{L}$, denoted by |-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{\mathsf{S}}_{I}$ is defined as follows:
@@
    The corresponding consequence operator is denoted by $W^{\mathsf{S}}_{I}$. The type-II $S$-entailment relation on $\mathscr{L}$, denoted by |-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{\mathsf{S}}_{II}$, is defined as follows.
    $\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{\mathsf{S}}_{II}\alpha\iff$ there $\text{there exists}~{}U\subseteq R~{}\text{with}~{}(\chi_{\Gamma},\chi_{W(\Gamma)})\in U~{}\text{such that}$
    for $\text{for all}~{}(v,w)\in U,v\models\Gamma~{}\text{implies that}~{}w\models\{\alpha\}$
@@
    Suszko and his followers admit as such an arbitrary family of arbitrary functions from the set of propositional formulas to the set of logical values. This last set can be identified with the set $2=\{0,1\}$, where 00 represents the value $\mathsf{false}$ and $1$ represents the value $\mathsf{true}$. Such a semantics can be called a Suszko $\ldots$
@@
    Let $\Gamma\subseteq\mathscr{L}$ such that $W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)\not\subseteq W^{\mathsf{S}}_{I}(\Gamma)$. Then there exists $\alpha\in W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$. So, there exists $(v,w)\in R$ such that $v\models\Gamma$, but $w\not\models\{\alpha\}$. Now, $v\models\Gamma$ implies that $\ast\models_{v}\Gamma$ and $w\not\models\{\alpha\}$ implies that ⊧̸$\ast\not\models_{w}\{\alpha\}∗ ⊧$. In other words, $\alpha\notin W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$- a contradiction. Thus $W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)\subseteq W^{\mathsf{S}}_{I}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ such that $W^{\mathsf{S}}_{I}(\Gamma)\not\subseteq W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$, Then there exists $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$. So, there exists $(\models_{v},\models_{w})\in S$ such that $\ast\models_{v}\Gamma$, but ⊧̸$\ast\not\models_{w}\{\alpha\}∗ ⊧$. Now, $\ast\models_{v}\Gamma$ implies that $v\models\Gamma$, and ⊧̸$\ast\not\models_{w}\{\alpha\}∗ ⊧$ implies that $w\not\models\{\alpha\}$. In other words, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$- a contradiction. Thus $W^{\mathsf{S}}_{I}(\Gamma)\subseteq W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Therefore, $W_{\mathfrak{S}_{\mathsf{S}}}=W^{\mathsf{S}}_{I}$. ∎
@@ 3. $\bullet$
[…]
@@ 2. $\bullet$
[…]
@@
[…]
@@
[…]
@@
    Let $(\mathscr{L},W)$ be a $wct$-type logical structure and $\mathsf{S}=(\mathbf{B},R_{wct},\models,\mathcal{P}(\mathscr{L}))$ be a normal $S$-semantics for $\mathscr{L}$ satisfying the following properties.
@@ 5 What is a Many-valued Logical Structure?
    [T]his characterization is purely matrical and there are several reasons to wish not to be blocked within matrix theory. Firstly, because the intuitive idea of many-valuedness does not necessarily depend on matrix theory, and secondly, because the rise of two-valued non-matrical semantics has shed a new light on the $\ldots$.
@@
    One cannot help but notice the similarity between these formulations of the principle of bivalence and $\mathbf{ST}$. Consider $\mathbf{PoB}^{P}_{1}$ for example. If one subscribes to the truth-theoretic conception of the notion of logical validity -- the view that in logical inferences, ‘truth’ is necessarily preserved from premise(s) to conclusion(s) -- and maintains that there is a generic notion of ‘truth’, then $\mathbf{ST}$ can be seen as a principle of bivalence of the form $\mathbf{PoB}^{P}_{1}$. To see this, suppose $P$ is a property and a logical structure $(\mathscr{L},W)$ satisfies $P$ iff it has an adequate bivalent semantics. This means that every element of $\mathscr{L}$ satisfies (semantic) property (not necessarily the same as $P$) or not. In the context of ST, it is ‘designation’, which Suszko identifies with ‘truth’. We quote in this context the following passage from [36, pp. 8--9] (emphasis ours) .
@@
    We note the following passage from [36, pp. 9--10] (emphasis ours).
@@
    Given a $\kappa$-valued semantics of order $\lambda$ for a family of sets $\{\mathscr{L}_{i}\}_{i\in\lambda}$, the entailment relation induced by $\mathfrak{S}$, denoted by |-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$ is defined as follows. For all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$,
    $\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}\alpha\iff$ for $~{}\text{for all}~{}((\models_{i},a_{i}),(\models_{j},a_{j}))\in S~{}\text{and}~{}(m_{i},m_{j})\in\mathbf{M}_{i}\times\mathbf{M}_{j},$
@@
      |-=|-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$.
@@
      For every $\mu$-valued semantics $\mathfrak{T}=(\{\mathbf{N}_{i}\}_{i\in\lambda},\{\models_{i}\}_{i\in\lambda},\{A_{i}\}_{i\in\lambda},S,\{\mathcal{P}(\mathscr{L}_{i})\}_{i\in\lambda})$ of order $\lambda$ for $\{\mathscr{L}_{i}\}_{i\in\lambda}$, with $\mu<\kappa$, |-|-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,\neq\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\mu,\lambda)}_{\mathfrak{S}}$.
@@
    It is imperative to emphasise that even though we defined of |-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$ in a specific manner, other definitions are possible. Thus, a definition of $\kappa$-valued logical structure depends on how |-$\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$ is defined.
@@
      To distinguish between the use and mention of a word or symbol, we sometimes put quotation marks around that symbol whenever we mention it. In the context of (formal) object language and (formal) metalanguage, this means that for every formula $\alpha$ of the object language, ‘$\alpha$’ is a constant symbol of the metalanguage (see, [13, Footnote 136] and [12, Section 6.2]). This gives rise to a natural injective function from the metalanguage to the object language under consideration. Property (i) of Definition 5.2 has been formulated with this issue in mind.
@@ 4. $\bullet$
      An explicit definition of ‘$\kappa$-valued logical structure’ (Definition 5.2) and generalised versions of $\mathbf{ST}$ are discussed.
@@ 1. $\bullet$
      Is it possible to obtain an adequate $S$-semantics for every non-monotonic logical structure, perhaps by changing the notion of $\mathsf{S}$-entailment?
@@ 2. $\bullet$
    3. $\bullet$
      Given a set $\mathscr{L}$, is it possible to characterise $\mathcal{F}_{\kappa}(\mathscr{L})$ for each cardinal $\kappa\geq 1$?

### v2

@@ 4 Suszko Reduction: A Generalised Perspective
    Given an $S$-semantics $\mathsf{S}=(\mathbf{B},R,\models,\mathcal{P}(\mathscr{L}))$, the type-I $S$-entailment relation on $\mathscr{L}$, denoted by $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{\mathsf{S}}_{I}$ is defined as follows:
@@
    The corresponding consequence operator is denoted by $W^{\mathsf{S}}_{I}$. The type-II $S$-entailment relation on $\mathscr{L}$, denoted by $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{\mathsf{S}}_{II}$, is defined as follows.
    $\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{\mathsf{S}}_{II}\alpha\iff$ $\text{there exists}~U\subseteq R~\text{with}~(\chi_{\Gamma},\chi_{W(\Gamma)})\in U~\text{such that}$
    $\text{for all}~(v,w)\in U,v\models\Gamma~\text{implies that}~w\models\{\alpha\}$
@@
[…]
@@
    Let $\Gamma\subseteq\mathscr{L}$ such that $W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)\not\subseteq W^{\mathsf{S}}_{I}(\Gamma)$. Then there exists $\alpha\in W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$ such that $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$. So, there exists $(v,w)\in R$ such that $v\models\Gamma$, but $w\not\models\{\alpha\}$. Now, $v\models\Gamma$ implies that $\ast\models_{v}\Gamma$ and $w\not\models\{\alpha\}$ implies that $\ast\not\models_{w}\{\alpha\}$. In other words, $\alpha\notin W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$ - a contradiction. Thus $W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)\subseteq W^{\mathsf{S}}_{I}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$.
    Next, let $\Gamma\subseteq\mathscr{L}$ such that $W^{\mathsf{S}}_{I}(\Gamma)\not\subseteq W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$, Then there exists $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$ such that $\alpha\notin W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$. So, there exists $(\models_{v},\models_{w})\in S$ such that $\ast\models_{v}\Gamma$, but $\ast\not\models_{w}\{\alpha\}$. Now, $\ast\models_{v}\Gamma$ implies that $v\models\Gamma$, and $\ast\not\models_{w}\{\alpha\}$ implies that $w\not\models\{\alpha\}$. In other words, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$ - a contradiction. Thus $W^{\mathsf{S}}_{I}(\Gamma)\subseteq W_{\mathfrak{S}_{\mathsf{S}}}(\Gamma)$ for all $\Gamma\subseteq\mathscr{L}$. Therefore, $W_{\mathfrak{S}_{\mathsf{S}}}=W^{\mathsf{S}}_{I}$. ∎
@@ 3. $\bullet$
    Now, let $\Gamma\in\mathsf{K}$, $\alpha\in\Gamma$ and $(v,w)\in R$ such that $v(\Gamma)\subseteq\{1\}$ and $w(\alpha)\neq 1$ (such a $(v,w)$ exists by hypothesis). Thus, $\alpha\notin W^{\mathsf{S}}_{I}(\Gamma)$, which implies that $\Gamma\subseteq\mathscr{L}\setminus W^{\mathsf{S}}_{I}(\Gamma)$. Hence, $\Gamma\subseteq\mathscr{L}\setminus W^{\mathsf{S}}_{I}(\Gamma)$ for all $\Gamma\in\mathsf{K}$. Thus, $W^{\mathsf{S}}_{I}$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$ and hence an $s^{\mathsf{K}}_{\kappa}$-consequence operator. ∎
@@ 2. $\bullet$
    $\Gamma\subseteq\mathscr{L}\setminus W(\Sigma)$ $\implies W(\Gamma)\subseteq W(\mathscr{L}\setminus W(\Sigma))$ $\text{(by monotonicity of}~W)$
@@
    $\implies\alpha\in W(\Sigma)$ $\text{(since}~\alpha\in W(\Gamma))$
@@
    Next, let $\Gamma\subseteq\mathscr{L}$ and $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$. If $\Gamma\in\mathsf{K}$ and $\mathscr{L}\setminus W(\Gamma)\in\mathsf{K}$, then $(\chi_{\mathscr{L}\setminus W(\Gamma)},\chi_{W(\Gamma)})\in R_{s^{\mathsf{K}}_{\kappa}}$. Since $W$ is $\text{anti-reflexive}^{\mathsf{K}}_{\kappa}$, and $\Gamma\in\mathsf{K}$, $\Gamma\subseteq\mathscr{L}\setminus W(\Gamma)$, which implies that $\chi_{\mathscr{L}\setminus W(\Gamma)}(\Gamma)\subseteq\{1\}$. Therefore, since $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$, we have $\chi_{W(\Gamma)}(\alpha)=1$, i.e., $\alpha\in W(\Gamma)$. If, on the other hand, $\Gamma\notin\mathsf{K}$ or $\mathscr{L}\setminus W(\Gamma)\notin\mathsf{K}$, then $(\chi_{\Gamma},\chi_{W(\Gamma)})\in R_{s^{\mathsf{K}}_{\kappa}}$. Since $\chi_{\Gamma}(\Gamma)\subseteq\{1\}$ and $\alpha\in W^{\mathsf{S}}_{I}(\Gamma)$, $\chi_{W(\Gamma)}(\alpha)=1$, i.e., $\alpha\in W(\Gamma)$.
@@
    Let $(\mathscr{L},W)$ be a wctwct-type logical structure and $\mathsf{S}=(\mathbf{B},R_{wct},\models,\mathcal{P}(\mathscr{L}))$ be a normal $S$-semantics for $\mathscr{L}$ satisfying the following properties.
@@ 5 What is a Many-valued Logical Structure?
[…]
@@
    One cannot help but notice the similarity between these formulations of the principle of bivalence and $\mathbf{ST}$. Consider $\mathbf{PoB}^{P}_{1}$ for example. If one subscribes to the truth-theoretic conception of the notion of logical validity - the view that in logical inferences, ‘truth’ is necessarily preserved from premise(s) to conclusion(s) - and maintains that there is a generic notion of ‘truth’, then $\mathbf{ST}$ can be seen as a principle of bivalence of the form $\mathbf{PoB}^{P}_{1}$. To see this, suppose $P$ is a property and a logical structure $(\mathscr{L},W)$ satisfies $P$ iff it has an adequate bivalent semantics. This means that every element of $\mathscr{L}$ satisfies (semantic) property (not necessarily the same as $P$) or not. In the context of ST, it is ‘designation’, which Suszko identifies with ‘truth’. We quote in this context the following passage from [36, pp. 8-9] (emphasis ours) .
@@
    We note the following passage from [36, pp. 9-10] (emphasis ours).
@@
    Given a $\kappa$-valued semantics of order $\lambda$ for a family of sets $\{\mathscr{L}_{i}\}_{i\in\lambda}$, the entailment relation induced by $\mathfrak{S}$, denoted by $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$ is defined as follows. For all $\Gamma\cup\{\alpha\}\subseteq\mathscr{L}$,
    $\Gamma\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}\alpha\iff$ $~\text{for all}~((\models_{i},a_{i}),(\models_{j},a_{j}))\in S~\text{and}~(m_{i},m_{j})\in\mathbf{M}_{i}\times\mathbf{M}_{j},$
@@
      $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,=\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$.
@@
      For every $\mu$-valued semantics $\mathfrak{T}=(\{\mathbf{N}_{i}\}_{i\in\lambda},\{\models_{i}\}_{i\in\lambda},\{A_{i}\}_{i\in\lambda},S,\{\mathcal{P}(\mathscr{L}_{i})\}_{i\in\lambda})$ of order $\lambda$ for $\{\mathscr{L}_{i}\}_{i\in\lambda}$, with $\mu<\kappa$, $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}\,\neq\,\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\mu,\lambda)}_{\mathfrak{S}}$.
@@
    It is imperative to emphasise that even though we defined of $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$ in a specific manner, other definitions are possible. Thus, a definition of $\kappa$-valued logical structure depends on how $\mathrel{|}\joinrel\mkern-0.5mu\mathrel{-}^{(\kappa,\lambda)}_{\mathfrak{S}}$ is defined.
@@
      To distinguish between the use and mention of a word or symbol, we sometimes put quotation marks around that symbol whenever we mention it. In the context of (formal) object language and (formal) metalanguage, this means that for every formula $\alpha$ of the object language, $\alpha$’ is a constant symbol of the metalanguage (see, [13, Footnote 136] and [12, Section 6.2]). This gives rise to a natural injective function from the metalanguage to the object language under consideration. Property (i) of Definition 5.2 has been formulated with this issue in mind.
@@ 4. $\bullet$
      An explicit definition of $\kappa$-valued logical structure’ (Definition 5.2) and generalised versions of $\mathbf{ST}$ are discussed.
@@ 1. $\bullet$
      Is it possible to obtain an adequate $S$-semantics for every non-monotonic logical structure, perhaps by changing the notion of $S$-entailment?